OpenStudy (anonymous):

Solve for x in the following proportion: 3 over 7 equals 2 over x A. x = 14 B. x = 7 C. x = 21/2 D. x = 14/3 this is d or A

OpenStudy (anonymous):

D

OpenStudy (whpalmer4):

Solve by cross-multiplication \[\frac{3}{7}=\frac{2}{x}\]\[3*x=2*7\] \[x=\]

OpenStudy (anonymous):

x = 14/3